### v1

@@ Abstract
Stochastic variance reduced gradient (SVRG) is an accelerated version of stochastic gradient descent based on variance reduction, and is promising for solving large-scale inverse problems. In this work, we analyze SVRG and a regularized version that incorporates a priori knowledge of the problem, for solving linear inverse problems in Hilbert spaces. We prove that, with suitable constant step size schedules and regularity conditions, the regularized SVRG can achieve optimal convergence rates in terms of the noise level without any early stopping rules, and standard SVRG is also optimal for problems with nonsmooth solutions under a priori stopping rules. The analysis is based on an explicit error recursion and suitable prior estimates on the inner loop updates with respect to the anchor point. Numerical experiments are provided to complement the theoretical analysis.
@@ 1 Introduction
with bounded linear operators $A_{{\dagger},i}:X\rightarrow Y_{i}$ between Hilbert spaces $X$ and $Y_{i}$ equipped with norms $\|\cdot\|_{X}$ and $\|\cdot\|_{Y_{i}}$, respectively, and the superscript $T$ denoting the vector transpose. $x\in X$ denotes the unknown signal of interest and $y_{\dagger}=(y_{{\dagger},1},\cdots,y_{{\dagger},n})^{T}\in Y$ denotes the exact data, i.e., $y_{\dagger}=A_{\dagger}x_{\dagger}$ with $x_{\dagger}$ being the minimum-norm solution relative to the initial guess $x_{0}$, cf. (2.1). In practice, we only have access to a noisy version $y^{\delta}$ of the exact data $y_{\dagger}$, given by
@@
where $\xi=(\xi_{1},\cdots,\xi_{n})^{T}\in Y$ is the noise in the data with a noise level $\delta=\|\xi\|_{Y}:=\sqrt{\sum_{i=1}^{n}\|\xi_{i}\|_{Y_{i}}^{2}}.$ Below we assume $\delta<1$. Linear inverse problems arise in many practical applications, e.g., computed tomography [9, 22, 4] and positron emission tomography [10, 16, 23].
Stochastic iterative algorithms, including stochastic gradient descent (SGD) [21, 11, 25] and stochastic variance reduced gradient (SVRG) [15, 24, 13], have gained much interest in the inverse problems community in recent years, due to their excellent scalability with respect to data size. We refer interested readers to the recent surveys [2, 12] for detailed discussions. Specifically, consider the following optimization problem
$J(x)=\tfrac{1}{2n}\|A_{\dagger}x-y^{\delta}\|_{Y}^{2}=\tfrac{1}{n}\sum_{i=1}^{n}f_{i}(x),\quad\mbox{with}\quad f_{i}(x)=\tfrac{1}{2}\|A_{{\dagger},i}x-y_{i}^{\delta}\|_{Y_{i}}^{2}.$
Given an initial guess $x_{0}^{\delta}\equiv x_{0}\in X$, SGD is given by
@@
while SVRG reads
$x_{k+1}^{\delta}=x_{k}^{\delta}-\eta_{k}\big(f^{\prime}_{i_{k}}(x_{k}^{\delta})-f^{\prime}_{i_{k}}(x_{[k/M]M}^{\delta})+J^{\prime}(x_{[k/M]M}^{\delta})\big),\quad k=0,1,\cdots,$
where $\{\eta_{k}\}_{k\geq 0}\subset(0,\infty)$ is the step size schedule, the index $i_{k}$ is sampled uniformly at random from the set $\{1,\ldots,n\}$, $M$ is the frequency of computing the full gradient, and $\cdot$] denotes taking the integral part of a real number.
By combining the full gradient $J^{\prime}(x_{[k/M]M}^{\delta})$ of the objective $J$ at the anchor point $x_{[k/M]M}^{\delta}$ with a random gradient gap $f_{i_{k}}^{\prime}(x_{k}^{\delta})-f_{i_{k}}^{\prime}(x_{[k/M]M}^{\delta})$, SVRG can accelerate the convergence of SGD and has become very popular in stochastic optimization [1, 5]. Its performance depends on the frequency $M$ of computing the full gradient, and $M$ was suggested to be $2n$ and $5n$ for convex and nonconvex optimization, respectively [15]. In practice, there are several variants of SVRG, depending on the choice of the anchor point, e.g., last iterate and randomly selected iterate within the inner loop. In this work, we focus on the version given in Algorithm 1, where $A_{\dagger}^{*}$ and $A_{{\dagger},i}^{*}$ denote the adjoints of the operators $A_{\dagger}$ and $A_{{\dagger},i}$, respectively.
@@
The low-rank nature of $A_{\dagger}$ implies that one can extract a low-rank subspace. Several works have proposed subspace / low-rank versions of stochastic algorithms [18, 6, 19, 20, 8]. Let $A:=(A_{1},\cdots,A_{n})^{T}$ approximate $A_{\dagger}$. Using $A$ in place of $A_{\dagger}$ in Algorithm 1 gives Algorithm 2, termed as regularized SVRG (rSVRG) below. rSVRG may be interpreted as integrating learned prior into SVRG, if $A$ is generated from paired training dataset $\{x^{(j)},y^{(j)}\}_{j=1}^{N}$. The regularization provided by the learned prior may relieve the need of early stopping.
@@
The mathematical theory of SVRG for inverse problems from the perspective of regularization theory has not been fully explored, and only recently has its convergence rate for solving linear inverse problems been investigated [13, 14]. In this work, we establish convergence of both rSVRG and SVRG for solving linear inverse problems. See Theorem 2.1 for convergence rates in terms of the iteration index $k$, Corollary 2.1 for convergence rates in terms of $\delta$, and Corollary 2.3 for regularizing property. Note that rSVRG has a built-in regularization mechanism without any need of early stopping rules and can outperform SVRG (i.e., higher accuracy), cf. Section 4. Moreover, we establish the (optimal) convergence rates in both expectation and uniform sense for both SVRG (when combined with a priori stopping rules) and rSVRG (cf. Theorem 2.1 and Corollary 2.1), while the prior works [13, 14] only studied convergence rates in expectation. For SVRG, the condition for its optimal convergence rate in expectation is more relaxed than that in [13]. However, unlike the results in [13], SVRG loses its optimality for smooth solutions under the relaxed condition. For the benchmark source condition $x_{\dagger}-x_{0}\in\mathrm{Range}(A_{\dagger}^{*})$ studied in [14], the condition is either comparable or more relaxed; see Remark 2.1 for the details.
The rest of the work is organized as follows. In Section 2, we present and discuss the main result, i.e., the convergence rate for (r)SVRG in Theorem 2.1. We present the proof in Section 3. Then in Section 4, we present several numerical experiments to complement the analysis, which indicate the advantages of rSVRG over standard SVRG and Landweber method. Finally, we conclude this work with further discussions in Section 5. In Appendix A, we collect lengthy and technical proofs of several technical results. Throughout, we suppress the subscripts in the norms and inner products, as the spaces are clear from the context.
@@ 2 Main result and discussions
and the operator $A$, for analyzing the convergence of the rSVRG. We denote the operator norm of $A_{i}$ by $\|A_{i}\|$ and that of $A$ by $\|A\|\leq\sqrt{\sum_{i=1}^{n}\|A_{i}\|^{2}}$. $\mathcal{N}(A_{\dagger})$ denotes the null space of $A_{\dagger}$.
@@
    There exist $\nu>0$ and $w\in\mathcal{N}(A_{\dagger})^{\perp}$ such that $x_{\dagger}-x_{0}=B_{\dagger}^{\nu}w$ and $\|w\|<\infty$, with $B_{\dagger}=n^{-1}A_{\dagger}^{*}A_{\dagger}$ and $\mathcal{N}(A_{\dagger})^{\perp}$ being the orthogonal complement of $\mathcal{N}(A_{\dagger})$.
@@
    Let $a\geq 0$ be a constant. When $a=0$, set $A=A_{\dagger}$. When $a>0$, let $A_{\dagger}$ be a compact operator with $\{\sigma_{j},\varphi_{j},\psi_{j}\}_{j=1}^{\infty}$ being its singular values and vectors, i.e., $A_{\dagger}(\cdot)=\sum_{j=1}^{\infty}\sigma_{j}\langle\varphi_{j},\cdot\rangle\psi_{j}$, such that $\{\sigma_{j}\}_{j=1}^{\infty}\subset[0,\infty)$, $\sigma_{j}\geq\sigma_{j^{\prime}}\geq a\delta^{b}>0$ for any $j\leq j^{\prime}\leq J$, and $\sigma_{j}<a\delta^{b}$ for any $j>J$, with some $b>0$. Set $A(\cdot)=\sum_{j=1}^{J}\sigma_{j}\langle\varphi_{j},\cdot\rangle\psi_{j}$.
The constant step size in Assumption 2.1(i) is commonly employed by SVRG [15]. (ii) is commonly known as the source condition [3], which imposes certain regularity on the initial error $x_{\dagger}-x_{0}$ and is crucial for deriving convergence rates for iterative methods. Without the condition, the convergence of regularization methods can be arbitrarily slow [3]. (iii) assumes that the operator $A$ captures important features of $A_{\dagger}$, and can be obtained by the truncated SVD of $A_{\dagger}$ that retains principal singular values $\sigma_{j}$ such that $\sigma_{j}\geq a\delta^{b}$. When $a=0$, $A=A_{\dagger}$ and rSVRG reduces to the standard SVRG.
Let $\mathcal{F}_{k}$ denote the filtration generated by the random indices $\{i_{0},i_{1},\ldots,i_{k-1}\}$, $\mathcal{F}={\bigvee_{k=1}^{\infty}}\mathcal{F}_{k}$, $(\Omega,\mathcal{F},\mathbb{P})$ denote the associated probability space, and $\mathbb{E}[\cdot]$ denote taking the expectation with respect to the filtration $\mathcal{F}$. The (r)SVRG iterate $x_{k}^{\delta}$ is random but measurable with respect to the filtration $\mathcal{F}_{k}$. Now, we state the main result on the error $e_{k}^{\delta}=x_{k}^{\delta}-x_{\dagger}$ of the (r)SVRG iterate $x_{k}^{\delta}$ with respect to $x_{\dagger}$. Below we follow the convention $k^{0}:=\ln k$, and let
$\overline{C_{0}}$ $:=\max\big(\|A\|^{-1}(5Ln^{-1}M)^{-\frac{1}{2}},L^{-1}(10+\ln M)^{-1}\big),$
$C_{0}$ $:=\big(100\sqrt{L}M\|A\|\ln(2e^{2}n\sqrt{L}\|A\|^{-1})\big)^{-1}.$
@@
The next corollary follows directly from Theorem 2.1.
@@
    When $a>0$ and $b=(1+2\nu)^{-1}$, i.e., rSVRG, for any small $\epsilon>0$,
@@
    When $a=0$, i.e., SVRG,
@@
With a suitable choice of $A$, rSVRG can achieve optimal convergence rates without any early stopping rule. SVRG is also optimal with a priori stopping rules for $\nu\in(0,\frac{1}{2})$. These rates are identical with that of SVRG in [13, 14]. Note that, when $\nu\in(0,\frac{1}{2}]$, the condition for optimal convergence rates in expectation of standard SVRG is more relaxed than that in [13], which requires also a special structure on $A_{\dagger}$ and the step size $c_{0}\leq\mathcal{O}\big((Mn^{-1}\|A\|^{2})^{-1}\big)$. It is comparable with that in [14] for small $M$ and more relaxed than that for relatively large $M$.
Assumption 2.1(iii) is to simplify the proof in Section 3. In fact, without (iii), the result of SVRG (i.e., $a=0$) in Theorem 2.1 holds trivially, while the result for rSVRG (i.e., $a>0$) remains valid when $A_{\dagger}$ can be approximated by some operator $A$ suitably; see the next corollary.
@@
In the absence of the source condition in Assumption 2.1(ii), the regularizing property of (r)SVRG remains valid in expectation and in the uniform sense.
@@
Let Assumption 2.1(i) and (iii) hold. Then rSVRG is regularizing itself, and SVRG is regularizing when equipped with a suitable a priori stopping rule.

### v2

@@ Abstract
Stochastic variance reduced gradient (SVRG) is an accelerated version of stochastic gradient descent based on variance reduction, and is promising for solving large-scale inverse problems. In this work, we analyze SVRG and a regularized version that incorporates a priori knowledge of the problem, for solving linear inverse problems in Hilbert spaces. We prove that, with suitable constant step size schedules and regularity conditions, the regularized SVRG can achieve optimal convergence rates in terms of the noise level without any early stopping rules, provided that the truncation level is chosen suitably, and standard SVRG is also optimal for problems with nonsmooth solutions under a priori stopping rules. The analysis is based on an explicit error recursion and suitable a priori estimates on the inner loop updates with respect to the anchor point. Numerical experiments are provided to complement the theoretical analysis.
@@ 1 Introduction
with bounded linear operators $A_{{\dagger},i}:X\rightarrow Y_{i}$ between Hilbert spaces $X$ and $Y_{i}$ equipped with norms $\|\cdot\|_{X}$ and $\|\cdot\|_{Y_{i}}$, respectively, and the superscript $T$ denoting the vector transpose. The element $x\in X$ denotes the unknown signal of interest and $y_{\dagger}=(y_{{\dagger},1},\cdots,y_{{\dagger},n})^{T}\in Y$ denotes the exact data, i.e., $y_{\dagger}=A_{\dagger}x_{\dagger}$ with $x_{\dagger}$ being the minimum-norm solution relative to the initial guess $x_{0}$, cf. (2.1) for the definition of $x_{\dagger}$. In practice, we only have access to a noisy version $y^{\delta}$ of the exact data $y_{\dagger}$, given by
@@
where $\xi=(\xi_{1},\cdots,\xi_{n})^{T}\in Y$ is the noise in the data with a noise level $\delta=\|\xi\|_{Y}:=\sqrt{\sum_{i=1}^{n}\|\xi_{i}\|_{Y_{i}}^{2}}.$ Below we assume $\delta<1$. Linear inverse problems arise in many practical applications, e.g., computed tomography [11, 30, 5], photoacoustic tomography [28, 32] and positron emission tomography [12, 22, 31].
Stochastic iterative algorithms, including stochastic gradient descent (SGD) [29, 14, 13, 16, 35] and stochastic variance reduced gradient (SVRG) [20, 34, 17], have gained much interest in the inverse problems community in recent years, due to their excellent scalability with respect to data size. We refer interested readers to the recent surveys [3, 15] for detailed discussions. Specifically, consider the following ill-posed optimization problem
$J(x)=\frac{1}{2n}\|A_{\dagger}x-y^{\delta}\|_{Y}^{2}=\frac{1}{n}\sum_{i=1}^{n}f_{i}(x),\quad\mbox{with}\quad f_{i}(x)=\frac{1}{2}\|A_{{\dagger},i}x-y_{i}^{\delta}\|_{Y_{i}}^{2}.$
Given an initial guess $x_{0}^{\delta}\equiv x_{0}\in X$, the SGD iteration is given by
@@
while the SVRG iteration reads
$x_{k+1}^{\delta}=x_{k}^{\delta}-\eta_{k}\big(f^{\prime}_{i_{k}}(x_{k}^{\delta})-f^{\prime}_{i_{k}}(x_{[\frac{k}{M}]M}^{\delta})+J^{\prime}(x_{[\frac{k}{M}]M}^{\delta})\big),\quad k=0,1,\cdots,$
where $\{\eta_{k}\}_{k\geq 0}\subset(0,\infty)$ is the step size schedule, the index $i_{k}$ is sampled uniformly at random from the index set $\{1,\ldots,n\}$, $M$ is the frequency of computing the full gradient, and the notation $\cdot$] denotes taking the integral part of a real number.
By combining the full gradient $J^{\prime}(x_{[\frac{k}{M}]M}^{\delta})$ of the objective $J$ at the anchor point $x_{[\frac{k}{M}]M}^{\delta}$ with a random gradient gap $f_{i_{k}}^{\prime}(x_{k}^{\delta})-f_{i_{k}}^{\prime}(x_{[\frac{k}{M}]M}^{\delta})$, SVRG can accelerate the convergence of SGD and has become very popular in stochastic optimization [2, 6]. Its performance depends on the frequency $M$ of computing the full gradient, and $M$ was suggested to be $2n$ and $5n$ for convex and nonconvex optimization, respectively [20]. In practice, there are several variants of SVRG, depending on the choice of the anchor point, e.g., last iterate and randomly selected iterate within the inner loop. In this work, we focus on the version given in Algorithm 1, where $A_{\dagger}^{*}$ and $A_{{\dagger},i}^{*}$ denote the adjoints of the operators $A_{\dagger}$ and $A_{{\dagger},i}$, respectively.
@@
The low-rank nature of $A_{\dagger}$ implies that one can extract a low-rank subspace. Several works have proposed subspace / low-rank versions of stochastic algorithms [24, 7, 25, 26, 10]. Let $A:=(A_{1},\cdots,A_{n})^{T}$ approximate $A_{\dagger}$. Using $A$ in place of $A_{\dagger}$ in Algorithm 1 gives Algorithm 2, termed as regularized SVRG (rSVRG) below. The version described in Algorithm 2 follows Algorithm 1 directly, and there are also other variants. One such variant based on an approximate singular value decomposition (SVD) $A(\cdot)=\sum_{j=1}^{J}\sigma_{j}(\varphi_{j},\cdot)\psi_{j}$ of $A_{\dagger}$ may take the form $A_{j}(\cdot)=\sigma_{j}(\varphi_{j},\cdot)\psi_{j}$, $j=1,\ldots,J$. This alternative version is denoted by rSVRG below. Numerically computing SVD directly can be expensive, although randomized SVD [8] may allow extracting low-rank approximations efficiently. See Section 4 for a comparison between rSVRG and rSVRG. Conceptually speaking, rSVRG may be interpreted as integrating a learned prior into SVRG, if the approximation $A$ is generated from a paired training dataset $\{x^{(j)},y^{(j)}\}_{j=1}^{N}$. The regularization provided by the learned prior may relieve the need of early stopping.
@@
The mathematical theory of SVRG for inverse problems from the perspective of regularization theory has not been fully explored, and only recently has its convergence rate for solving linear inverse problems been investigated [17, 19]. In this work, we establish the convergence of both rSVRG and SVRG for solving linear inverse problems. See Theorem 2.1 for the error bounds in terms of the iteration index $k$, Corollary 2.1 for the convergence rates in terms of $\delta$, and Corollary 2.3 for regularizing property. Note that rSVRG inherits the regularization mechanism of the truncated SVD, without any need of early stopping rules, provided that the truncation level is chosen properly according to the noise level, and empirically, it can outperform SVRG (i.e., higher accuracy), cf. the numerical results in Section 4. Moreover, we establish the (optimal) convergence rates in both expectation and uniform sense for both SVRG (when combined with a priori stopping rules) and rSVRG (cf. Theorem 2.1 and Corollary 2.1), while the prior works [17, 19] only studied convergence rates in expectation. Note that the work [17] is based on the spectral analysis of bounded linear operators, whereas the work [19] is based on a Lyapunov type analysis. In this work, we follow the framework of the spectral analysis in [17]. For SVRG, the condition for its optimal convergence rate in expectation is more relaxed than that in [17]. However, unlike the results in [17], SVRG loses its optimality for smooth solutions under the relaxed condition. For the benchmark source condition $x_{\dagger}-x_{0}\in\mathrm{Range}(A_{\dagger}^{*})$ studied in [19], the condition is either comparable or more relaxed; see Remark 2.1 for further details.
The rest of the work is organized as follows. In Section 2, we present and discuss the main results, i.e., the convergence rate for (r)SVRG in Theorem 2.1 and the regularizing properties in Corollary 2.3. We present the detailed proofs in Section 3. Then in Section 4, we present several numerical experiments to complement the analysis. Finally, we conclude this work with further discussions in Section 5. In Appendix A, we collect lengthy and technical proofs of several technical results. Throughout, we suppress the subscripts in the norms and inner products, as the spaces are clear from the context.
@@ 2 Main result and discussions
and the operator $A$, for analyzing the convergence of rSVRG. We denote the operator norm of $A_{i}$ by $\|A_{i}\|$ and that of $A$ by $\|A\|\leq\sqrt{\sum_{i=1}^{n}\|A_{i}\|^{2}}$. $\mathcal{N}(A_{\dagger})$ denotes the null space of $A_{\dagger}$.
@@
    There exist $\nu>0$ and $w\in\mathcal{N}(A_{\dagger})^{\perp}$ such that $x_{\dagger}-x_{0}=B_{\dagger}^{\nu}w$, with $B_{\dagger}=n^{-1}A_{\dagger}^{*}A_{\dagger}$ and $\mathcal{N}(A_{\dagger})^{\perp}$ being the orthogonal complement of $\mathcal{N}(A_{\dagger})$.
@@
    Let $a\geq 0$ be a constant. When $a=0$, set $A=A_{\dagger}$. When $a>0$, let $A_{\dagger}$ be a compact operator with $\{\sigma_{j},\varphi_{j},\psi_{j}\}_{j=1}^{\infty}$ being its singular values and functions, i.e., $A_{\dagger}(\cdot)=\sum_{j=1}^{\infty}\sigma_{j}\langle\varphi_{j},\cdot\rangle\psi_{j}$, such that $\{\sigma_{j}\}_{j=1}^{\infty}\subset[0,\infty)$, $\sigma_{j}\geq\sigma_{j^{\prime}}\geq a\delta^{b}>0$ for any $j\leq j^{\prime}\leq J$, and $\sigma_{j}<a\delta^{b}$ for any $j>J$, with some $b>0$. Set $A(\cdot)=\sum_{j=1}^{J}\sigma_{j}\langle\varphi_{j},\cdot\rangle\psi_{j}$.
The constant step size in Assumption 2.1(i) is commonly employed by SVRG [20]. (ii) is commonly known as the source condition [4], which imposes certain regularity on the initial error $x_{\dagger}-x_{0}$ and is crucial for deriving convergence rates for (iterative) regularization methods. Without the condition, the convergence of regularization methods can be arbitrarily slow [4]. (iii) assumes that the operator $A$ captures the important features of $A_{\dagger}$, and can be obtained by the truncated SVD of $A_{\dagger}$ that retains principal singular values $\sigma_{j}$ such that $\sigma_{j}\geq a\delta^{b}$. When $a=0$, $A=A_{\dagger}$ and rSVRG reduces to the standard SVRG. (iii) is adopted to simplify the error analysis, and one can obtain similar results even if it is relaxed, which can be fulfilled by several data-driven models, including data-driven reduced order models [33, 27], autoencoder neural networks [21], and neural networks combined with model reduction [1, 18]. See Corollary 2.2 for details of the relevant convergence results in the absence of (iii).
Let $\mathcal{F}_{k}$ denote the filtration generated by the random indices $\{i_{0},i_{1},\ldots,i_{k-1}\}$, $\mathcal{F}={\bigvee_{k=1}^{\infty}}\mathcal{F}_{k}$, $(\Omega,\mathcal{F},\mathbb{P})$ denote the associated probability space, and $\mathbb{E}[\cdot]$ denote taking the expectation with respect to the filtration $\mathcal{F}$. The (r)SVRG iterate $x_{k}^{\delta}$ is random but measurable with respect to the filtration $\mathcal{F}_{k}$. Now, we state the main result on the error $e_{k}^{\delta}=x_{k}^{\delta}-x_{\dagger}$ of the (r)SVRG iterate $x_{k}^{\delta}$ with respect to the minimum-norm solution $x_{\dagger}$. Below we adopt the convention $k^{0}:=\ln k$, and let
$\overline{C_{0}}$ $:=\max\big(3^{-1}\|A\|^{-1}(5Ln^{-1}M)^{-\frac{1}{2}},9^{-1}L^{-1}(15+7\ln M)^{-1}\big),$
$C_{0}$ $:=\big(14^{2}\sqrt{L}M\|A\|\ln(2e^{2}n\sqrt{L}\|A\|^{-1})\big)^{-1}.$
Note that the constants $C_{0}$ and $\overline{C_{0}}$ are generally much smaller than $L^{-1}$, when compared with Assumption 2.1(i).
@@
The next corollary follows directly from Theorem 2.1. The shorthand notation $k(\delta)=\mathcal{O}(\delta^{-\frac{2}{1+2\nu}})$ means that there exist constants $0<C_{1}<C_{2}$ and some $\delta_{0}>0$, such that for any $\delta\in(0,\delta_{0}]$, there holds $C_{1}\delta^{-\frac{2}{1+2\nu}}\leq k(\delta)\leq C_{2}\delta^{-\frac{2}{1+2\nu}}.$
@@
    When $a>0$ and $b=(1+2\nu)^{-1}$, i.e., rSVRG, there holds for any small $\epsilon>0$,
@@
    When $a=0$, i.e., SVRG, there holds
@@
Note that the convergence rate $\mathcal{O}(\delta^{\frac{2\nu}{1+2\nu}})$ stated in Corollary 2.1 is identical with that for the Landweber method [4], which is known to be order-optimal. With a suitable choice of $A$, rSVRG can achieve optimal convergence rates without any early stopping rule, instead inheriting the regularizing property of the spectral cutoff. SVRG is also optimal with a priori stopping rules for $\nu\in(0,\frac{1}{2})$. These rates are identical with that of SVRG in [17, 19]. Note that, when $\nu\in(0,\frac{1}{2}]$, the condition for optimal convergence rates in expectation of standard SVRG is more relaxed than that in [17], which requires also a special structure on $A_{\dagger}$ and the step size $c_{0}\leq\mathcal{O}\big((Mn^{-1}\|A\|^{2})^{-1}\big)$. It is comparable with that in [19] for small $M$ and more relaxed than that for relatively large $M$.
Remark 2.2.
Note that the error estimates in Corollary 2.1 are derived under the assumption that the smoothness parameter $\nu$ is known a priori for determining the stopping index $k(\delta)$ for SVRG and the truncation level $J$ in rSVRG. However, the parameter $\nu$ is generally unknown in practical applications. Thus the a priori choice represents one significant restriction of the applicability of rSVRG. It is of much interest to develop a posteriori or data-driven choices of the truncation level $J$.
Assumption 2.1(iii) is to simplify the analysis of rSVRG in Section 3. The result for rSVRG (i.e., $a>0$) remains valid when $A_{\dagger}$ is approximated by some operator $A$ suitably; see the next corollary.
@@
In the absence of the source condition in Assumption 2.1(ii), the regularizing property of (r)SVRG remains valid both in expectation and in the uniform sense.
@@
Let Assumption 2.1(i) and (iii) hold. Then rSVRG is regularizing ((in expectation when $c_{0}<\bar{C_{0}}$ and in the uniform sense when $c_{0}<C_{0})$ when $b\in(0,1)$, and SVRG is regularizing ((in expectation when $c_{0}<\bar{C_{0}}$ and in the uniform sense when $c_{0}<C_{0})$ when the stopping index $k(\delta)$ satisfies the following conditions
$\lim_{\delta\to 0^{+}}k(\delta)=\infty\quad\mbox{and}\quad\lim_{\delta\to 0^{+}}\sqrt{k(\delta)}\delta=0.$